# Examining Differential Item Functioning (DIF) in Self-Reported Health Survey Data: Via Multilevel Modeling

### v1

@@ 4.1 Data preparation
The original NSDUH data file contains 2,605 variables and 59,069 respondents, composed by 27,047 males and 32,022 females aged 12 and older. We keep the following four variables for DIF analysis:
@@
    The item response variable (named “ADDPREV” in the survey). It contains the following responses to the selected survey item. It is recoded to have 1 for the response “Yes” and 0 for “No.”
@@
    The group membership variable (named “PDEN10”). It indicates the population density of the area where respondents are residing. It is recoded to have 1 for “high-density” (i.e., with one million or more people) and 0 for “low-density” (i.e., with fewer than one million people), based on Core Based Statistical Area (CBSA) classifications. These areas are also known as “large metropolitan” and “small metropolitan” areas [41].
@@
Appendix C displays information from the NSDUH codebook about how the selected categorical variables were originally coded.
Table 1: Sample size for each data cluster.
Level-2 clusters Group Response
Edu. level Low-density High-density No Yes
1 126 140 224 42
2 75 115 172 18
3 45 49 76 18
4 158 102 203 57
5 314 206 391 129
6 407 273 488 192
7 1228 1059 1590 697
8 6488 4430 7435 3483
9 5208 3895 5323 3780
10 2286 1563 2463 1386
11 7630 7874 9780 5724
Considering that the multilevel LR and the traditional LR use the binary logistic regression that are limited to the binary response data analysis, we focus on the respondents associated with “Yes” and “No,” removing the respondents that answer “Don’t know,” refuse to answer or leave a blank, or are led to skip the selected survey item due to the survey logic. So, for further analysis, we first drop 13,242 respondents associated with neither “Yes” nor “No.” Then, because we focus on DIF associated with population density, we exclude 2,156 respondents associated with non-CBSA classifications for which there is a lack of information of the population density. In this way, we retain 43,671 eligible respondents. They comprise 19,348 males and 24,323 females, with no missing values. Table 1 displays sample sizes for the level-2 clusters, which are large enough for multilevel modeling [42, 43].
@@ 4.2 Analysis and Results
We carried out DIF analyses using a series of models: three multilevel models associated with the multilevel LR shown in (1), and the single-level traditional LR model shown in (3). The first multilevel model we utilize does not contain explanatory variables in the level-1 submodel, which makes it the commonly known null model or empty model that sets a baseline for understanding the proportion of variance explained by the hierarchical structure of the data [27]. The second multilevel model is the exact model used in the multilevel LR, incorporating both the latent trait and group membership variables in the level-1 submodel. The third multilevel model adds the interaction term between the latent trait and group membership, the same as the interaction term in (3), to check if the model performance would be improved when adding this interaction. These three multilevel models are named “Model 0,” “Model 1,” and “Model 2.” The single-level model is labeled with “Model 3.”
The modeling results in Table 2 reveal several key insights regarding the performance of different models on DIF associated with population density in responses to the selected NSDUH item about depression. The null model (Model 0) serves as a baseline, showing the proportion of variance attributable to the hierarchical structure of the data. Its adjusted intraclass correlation coefficient [ICC; 24] of 0.071 indicates that 7.1% of the response variability stemmed from differences between education levels. This highlights the importance of considering this structure in DIF analyses with this data. Model 1, incorporating both the latent trait and group membership variables, displays a significant negative effect of high population density on the probability of responding “Yes” to the item. This suggests that this item has DIF, as individuals residing in high-density areas were less likely to provide the targeted response to the selected item compared to those in low-density areas, after accounting for all these individuals’ latent trait about depression. Model 2, which includes the interaction term, does not show a better model fit compared to Model 1 in this specific example. This implies that the interaction term does not substantially affect the model’s performance. Model 3, the single-level model, does not reveal a significant effect of high population density, which leads to the conclusion that there is no DIF, contrasting with the result from Model 1. This difference underscores the importance of model selection when conducting DIF analyses.
Overall, Models 1-2 outperform others, based on fit statistics. Both Model 1 and Model 2 show a low intercept variance ($\tau_{0}^{2}$ = 0.22), indicating an advantage in capturing more data variability. Additionally, they share the lowest AIC, BIC, and -2LL values, and the highest $R^{2}$ values. Their $R^{2}$ values signify that each of the two models explains 37.4% of the data variability through random and fixed effects and 33.1% solely through fixed effects. In contrast, the single-level model, Model 3, explains 26.9% of the variance, which is less favorable.
Table 2: Results from a series of models.
@@
$\sigma^{2}$ 3.29 3.29 3.29
N 43671 43671 43671 43671
AIC 56460 44197 44198 44697
BIC 56477 44232 44241 44732
-2LL 56455 44189 44188 44689
ICC 0.071 0.064 0.064
$R^{2}_{marg}$ 0.000 0.331 0.331 0.269
$R^{2}_{cond}$ 0.071 0.374 0.374
  • ${}^{*}p<0.05$; ${}^{**}p<0.01$; ${}^{***}p<0.001$.
Researchers might consider Model 2 when separate interpretations of uniform and nonuniform DIF are desired, analogous to the traditional LR approach. A significant interaction term in Model 2 suggests the presence of non-uniform DIF, where the effect of population density on the probability of responding “Yes” varies across the continuum of psychological distress scores. In contrast, a non-significant interaction term but a significant coefficient for the group membership variable would indicate uniform DIF, suggesting a constant difference in the probability of responding “Yes” between the high- and low-density groups regardless of the level of psychological distress.
It is important to acknowledge potential limitations of this analysis. While the sample size for each level-2 cluster appears sufficient based on existing recommendations, further exploration of the data distribution within each cluster could be beneficial [27]. This might involve examining potential skewness or outliers that could influence model estimates. Additionally, although we controlled for a latent trait variable reflecting psychological distress, the inclusion of other relevant covariates (e.g., socioeconomic status, access to healthcare) might provide a more comprehensive picture of potential DIF effects.
In conclusion, our results highlight the strengths of multilevel modeling for DIF analysis in health research with clustered data. By accounting for the hierarchical structure of the data (respondents nested within education levels), multilevel models provided a more accurate picture of the underlying relationships. This is evident in the superior model fit statistics (lower AIC, BIC, and -2LL values, and higher R-squared values) of multilevel models and the identification of DIF that was missed by the single-level model. Future research could explore the generalizability of our findings by applying multilevel DIF analysis to other self-reported health survey data.
@@ 5 Discussion
Given that single-level LR is not suitable for polytomous items or multiple groups and that we want to have a comparison between single-level and multilevel LR models, our empirical analysis centers on a dichotomous item and a variable involving two groups for DIF assessment. Rather than making a blanket recommendation in favor of multilevel models, we intend to raise awareness of how to use multilevel LR to analyze DIF and how to choose the appropriate model when a series of competing models are available. Different from some other studies that evaluated models via Monte Carlo simulation [e.g., 44], our empirical analysis examines a series of competing models based on model fit indices. Our findings indicate that multilevel modeling more effectively captures variability in our data compared to single-level modeling, which is crucial to accurate DIF detection.
Future research is crucial to comprehensively explore the cost-benefit balance of multilevel modeling for DIF analysis in dynamic health survey settings. Monte Carlo simulations can be a powerful tool to systematically evaluate model performance under a wide range of data structures. These simulations can vary factors such as ICC, number and size of clusters, and the magnitude and proportion of DIF items. Studies by Moineddin et al. [45] and [46] provide valuable groundwork, but further exploration is necessary. While initial efforts have been made to explore multilevel DIF analysis [e.g., 47, 48, 49], limitations exist regarding model construction, ICC calculation, and generalizability [23]. Scrutinizing model performance under various realistic scenarios that mimic real-world health survey data is essential. This includes exploring the impact of factors like missing data patterns (missing completely at random, missing at random, or missing not at random) and different types of DIF (uniform vs. non-uniform) on model performance. Also, researchers might consider developing or applying more advanced models capable of multilevel modeling for DIF analysis with polytomous items or more than two groups, ideally built on existing models without IRT-related assumptions such as ordinal logistic regression [50, 51] or nonparametric measures [e.g., 52]. By addressing these limitations, we can solidify the role of multilevel modeling in DIF analysis for health research. This will ultimately lead to more robust and reliable estimates captured through self-reported surveys.
@@
Appendix A R code
A.1 For data preparation
1# initialize
2load("NSDUH_2022.Rdata")
3d <- NSDUH_2022
4
5# keep relevant data
6d <- d[, c("addprev", # item
7 "PDEN10", # group membership
8 "KSSLR6MON", # latent trait
9 "IREDUHIGHST2" # Level-2 variable
10 )]
11
12# rename variables
13names(d) <- c("item", "group", "score", "level2")
14
15# remove invalid values
16d <- d[which(d\$group != 3), ] # non-CBSA areas
17d <- d[which(d\$item %inc(85, 94, 97, 98, 99) == FALSE),] # invalid response values
18
19# recode two variables
20d\$item[d\$item == 2] <- 0 # to have 1=Yes, 0=No
21d\$group[d\$group == 2] <- 0 # 1=high, 0=low
22
23# check missing values
24apply(d, MARGIN = 2, function(x){
25 table(x, useNA = "always")
26})
27
28# check level-2 sample sizes
29table(d\$level2, d\$group)
30table(d\$level2, d\$item)
31
32# finalize the data
33dta <- d
34dta\$group <- as.factor(dta\$group) # to make sure it is formated as a categorical variable, allowing glmer() to proceed with dummy coding
A.2 For DIF analyses
1library(lme4)
2
3# Model 0
4model0 <- glmer(item ~ 1 + (1 | level2), data = dta, family = binomial)
5summary(model0)
6tau00 <- as.data.frame(VarCorr(model0))[4]
7icc <- performance::icc(model0)
8resid <- tau00/icc\$ICC_adjusted - tau00
9#// residual variance
10
11# Model 1
12model1 <- glmer(item ~ score + group + (1 | level2), data = dta, family = binomial)
13summary(model1)
14tau00 <- as.data.frame(VarCorr(model1))[4]
15icc <- performance::icc(model1)
16resid <- tau00/icc\$ICC_adjusted - tau00
17
18# Model 2
19model2 <- glmer(item ~ score + group + score*group + (1 | level2), data = dta, family = binomial)
20summary(model2)
21tau00 <- as.data.frame(VarCorr(model2))[4]
22icc <- performance::icc(model2)
23resid <- tau00/icc\$ICC_adjusted - tau00
24
25# Model 3
26model3 <- glm(item ~ score + group + score*group, data = dta, family = "binomial")
27summary(model3)
28
29# compile results
30library(texreg)
31screenreg( list(model0, model1, model2, model3) )
32
33# obtain other statistics with this one
34sjPlot::tab_model(list (model0, model1, model2, model3))
Appendix B DIF Methods on Single-level Models
An extensive review of DIF methods can be found in Chen [23]. Table 3 summarizes the limitations of some popular DIF methods based on single-level models. In this appendix, we detail the Mantel-Haenszel procedure [MH; 17] and the Lord’s Wald test with worked examples, including the R code we used for computations. Our analyses were done via difR package [53] with R4.4.0 in RStudio.
Table 3: Limitations of the single-level DIF methods reviewed in Chen [23].
Method Limitations
MH [17] Limited to two groups; using observed, discrete total scores; assuming that scores are equally spaced, and that the common odds ratio remains constant.
LR [18] Limited to two groups; lacking a valid effect size measure; inflated errors with uniform DIF detection, with 3PL items.
Lord’s Wald test [19] Limited to two groups; item parameter estimates depend on the item set; lacking a valid effect size measure.
IRT-LR [20] Limited to two groups; computationally intensive; lacking a valid effect size measure.
B.1 The Mantel-Haenszel Procedure
The Mantel-Haenszel (MH) procedure, introduced by Holland and Thayer [17], is an extension of Pearson’s chi-square test of independence that uses odds ratios from tables to quantify the association between two dichotomous variables, conditional on a third variable that could be a confounder. In Holland’s work, a 2$\times$2 table is created to cross-classify the type of responses (i.e., correct vs. incorrect) by group (i.e., reference group vs. focal group) for the $k$th total score ($k$ = 1, 2, $\cdots$, $K$), demonstrated in Table 4. The row margins are fixed to 1 by design because an item response is either the targeted response (e.g., “Yes”) or not (e.g., “No”), and hence the cell values follow a binominal distribution for each row. The probabilities of answering an item correctly in the reference and focal groups are $\pi_{Rk}$ and $\pi_{Fk}$, respectively, conditional on the $k$th total score. Correspondingly, the probabilities of obtaining a targeted response in the reference and focal groups are $(1-\pi_{Rk})$ and $(1-\pi_{Fk})$, respectively.
Table 4: A 2$\times$2 table in probabilities (by group membership and item response) for the $k$th total score.
Yes (1) No (0) Total
Reference group $\pi_{Rk}$ $1-\pi_{Rk}$ 1
Focal group $\pi_{Fk}$ $1-\pi_{Fk}$ 1
In the MH, the null hypothesis is that there is no DIF, or in other words, that there is conditional independence between an examinee’s group membership and their response after accounting for their total scores. It is expressed as
$H_{0}\text{: }\dfrac{\pi_{Rk}}{1-\pi_{Rk}}=\dfrac{\pi_{Fk}}{1-\pi_{Fk}}\text{ for all }k.$
The alternative hypothesis is
$H_{1}\text{: }\dfrac{\pi_{Rk}}{1-\pi_{Rk}}=\alpha\dfrac{\pi_{Fk}}{1-\pi_{Fk}}\text{ for all }k,$
where $\alpha$ is the common odds ratio shared by the $k$ tables.
Probabilities in Table 4 can be expressed as counts, as shown in Table 5.
Table 5: A $2\times 2$ table in counts (by group and response) for the $k$th total score.
Yes (1) No (0)
Reference group $A_{k}$ $B_{k}$
Focal group $C_{k}$ $D_{k}$
Then the MH test statistic (with continuity correction to improve the approximations of significance levels) can be expressed as
$\chi^{2}_{MH}=\dfrac{(|\sum_{k}A_{k}-\sum_{k}E(A_{k})|-\dfrac{1}{2})^{2}}{\sum_{k}\mbox{var}(A_{k})},$
which follows a chi-square distribution, which assumes independence of observations.
Holland and Thayer [17] proposed an effect size measure of DIF, specified as
$\hat{\Delta}_{MH}=-2.35\ln(\hat{\alpha}),$
where the logarithm puts $\hat{\alpha}$ on an additively symmetric scale, making it convenient for interpretation. Since the 1990s, ETS has been flagging DIF items with $A$ (i.e., negligible/nonsignificant DIF), $B$ (i.e., slight to moderate DIF), and $C$ (i.e., moderate to large DIF) using $\chi^{2}_{MH}$ and $\hat{\Delta}_{MH}$, as documented in Zwick [54].
The MH method is simple and widely used in educational assessment development, but it has some limitations. Firstly, MH uses observed, discrete total scores as a proxy for a latent trait, resulting in information loss. Secondly, using discrete total scores as the third dimension for analyzing two-way tables involves unrealistic assumptions: that total scores are equally spaced and that the common odds ratio $\alpha$ remains constant for different total scores [54].
Application in R. Using the data of the respondents answering either “Yes” or “No” to the selected item we described in the main text, we found the MH test statistic for the item is 15.406, its $p$-value is 0.0001, which is highly significant, suggesting strong evidence against the null hypothesis that there is no DIF. The common odds ratio is greater than 1 ($\alpha_{MH}=1.0980$), suggesting that the respondents in the reference group is more likely to answer “Yes” to the selected item than the focal group.
However, it is important to consider both the statistical significance ($p$-value) and the effect size when interpreting these results, as a statistically significant finding may not always have practical significance. The effect size measure of DIF indicates a negligible effect ($\hat{\Delta}_{MH}=-0.2197$, labeled with $A$).
Table 6: Results from the Mantel-Haenszel procedure.
Measure Statistic $p$-value
$\chi^{2}_{MH}$ 15.406 ∗∗∗ 0.0001
$\alpha_{MH}$ 1.098
$\hat{\Delta}_{MH}$0.220
Effect Size $A$
  • ${}^{*}p<0.05$; ${}^{**}p<0.01$; ${}^{***}p<0.001$.
In summary, the MH test results indicate that DIF exists for the item. Nevertheless, its effect size is negligible, suggesting that the practical significance of this DIF might be small. The following R code demonstrates how to use an existing R function difMH() to do the MH analysis, as well as how to directly compute the MH test statistic without that function. We found that the MH test statistic is 15.406 via either of the procedures.
1# label the groups with "F" for reference groups and "R" for focal groups
2dta\$group <- ifelse(dta\$group == 1, "F", "R")
3
4# MH test (via an existing R function)
5library(difR)
6mh_result <- difMH(Data = as.data.frame(dta\$item), group = dta\$group, focal.name = ’F’, match = dta\$score)
7mh_result
8#// display results
9
10# MH test (via direct computation)
11l <- split(dta, as.factor(dta\$score))
12
13vals <- list()
14for(i in 1:length(l)){
15
16 x <- l[[i]]
17
18 d <- data.frame(table(x\$group, x\$item))
19
20 a <- d\$Freq[which(d\$Var1 == "R" & d\$Var2 == 1)]
21 n <- sum(d\$Freq)
22
23 m1 <- sum(d\$Freq[which(d\$Var1 == "R")])
24 n1 <- sum(d\$Freq[which(d\$Var2 == 1)])
25 a_expected <- m1*n1/n
26
27 m2 <- sum(d\$Freq[which(d\$Var1 == "F")])
28 n2 <- sum(d\$Freq[which(d\$Var2 == 0)])
29 variance <- (m1/10*m2/10*n1/10*n2/10)/(n^2*(n-1))*10^4
30 #// function: m1*m2*n1*n2/(n^2*(n-1)) -- we added "/10" and "*10^4" here to avoid the overflown error in R
31
32 t <- data.frame(cbind(a, a_expected, variance))
33 names(t) <- c("a", "a_expected", "variance")
34 vals[[i]] <- t
35}
36vals <- do.call(rbind, vals)
37
38chi <- (abs(sum(vals\$a) - sum(vals\$a_expected)) - 1/2)^2 / sum(vals\$variance)
39chi
40#// display the test statistic
B.2 Lord’s Wald Test
Lord’s Wald test [19], derived from the traditional Wald test [55], is historically significant. It requires the item response theory models (IRT) to estimate item parameters before computing DIF. Its two-sided null hypothesis is
$H_{0}\text{: }\beta-\beta_{0}=0,$ (4)
where $\beta$ denotes a population parameter such as a linear effect of an explanatory variable about the group membership in a model. The test statistic, in the form of a $z$ score that approximately follows a standard normal distribution, is
$z=\dfrac{\beta-\beta_{0}}{SE},$ (5)
where $SE$ is the standard error.
Specifically, Lord’s Wald test statistic is derived from the IRT-based item difficulty parameter from two groups for a studied item $i$, written as Langer [35]
$\hat{B}=\dfrac{\hat{b}_{F}-\hat{b}_{R}}{\sqrt{\mbox{var}(\hat{b}_{F})+\mbox{var}(\hat{b}_{R})}},$ (6)
where
$\hat{b}_{F}$ is the maximum likelihood estimate of the item difficulty parameter in the focal group;
$\hat{b}_{R}$ is the maximum likelihood estimate of the item difficulty parameter in the reference group;
$\mbox{var}(\hat{b}_{F})$ and $\mbox{var}(\hat{b}_{R})$ denote the sample variances of $\hat{b}_{F}$ and $\hat{b}_{R}$.
(6) is a special materialization of (5) for deriving the $z$ score, where the numerator is the tested parameter value and the denominator is the true standard error $SE$.
Lord extended (6) to include the IRT-based item discrimination and pseudo-guessing parameters for two groups. The hypothesis about the joint between-group difference derived from multiple IRT parameters is tested using a chi-square statistic expressed as
$\chi^{2}_{Lord}=\boldsymbol{\hat{\gamma}^{\prime}}\boldsymbol{\Sigma}^{-1}\boldsymbol{\hat{\gamma}},$
where
$\boldsymbol{\hat{\gamma}^{\prime}}$ is $\begin{bmatrix}\hat{a}_{F}-\hat{a}_{R}&\hat{b}_{F}-\hat{b}_{R}\end{bmatrix}$ when considering item discrimination together with item difficulty;
it is $\begin{bmatrix}\hat{a}_{F}-\hat{a}_{R}&\hat{c}_{F}-\hat{c}_{R}\end{bmatrix}$ when considering item discrimination together with item pseudo-guessing;
$\boldsymbol{\Sigma}^{-1}$ is the variance-covariance matrix of $\hat{\gamma}$.
Although Lord’s Wald test has not often been used for DIF detection in psychometric practice, it has been revived in modern DIF method discussions by Kim et al. [56], Langer [35], and Woods et al. [46]. Kim et al. [56] extended the method for DIF analyses with three or more groups. Langer [35] argued that the poor performance of Lord’s Wald test was partly due to its original item parameter estimation method. Woods et al. [46] compared improved versions of Lord’s Wald test.
As mentioned earlier, we first need to estimate the item parameters of this selected item when running Lord’s Wald test. It is important to highlight that estimating the parameters of a single item is not feasible because the corresponding IRT model would be under-identified. Moreover, the parameter estimate for an item fluctuates when varying the other items included in the analysis, owing to variations in the latent trait scale associated with each different item set. In practice, only one item set is needed to make conclusions about the DIF presence; this item set is chosen because it is deemed most appropriate to help describe the respondents.
Application in R. For this analysis, we used the Verbal Aggression data [57], made of the responses from 316 respondents to a questionnaire about verbal aggression. The response data were coded with 1 to mean that the respondent would want to respond to a frustrating situation aggressively, and with 0 to mean otherwise. To show the impact of the item set on Lord’s Wald test statistic, we explored this DIF analysis of our targeted item named “S1wantCurse” when using four different item sets that involved this item in addition to one, two, three other items, and all the other items in the Verbal data set, respectively. All these items were calibrated using the one-parameter item response model for parsimony. The selected items are specified below, for each of the item sets labeled with A, B, C, and D.
Table 7: Selected items in the three item sets.
Item Set N. of Items Item Names
A 2 S1wantCurse, S1WantScold.
B 3 S1wantCurse, S1WantScold,
S1WantShout.
C 4 S1wantCurse, S1WantScold,
S1WantShout, S2WantCurse.
D 24 S1wantCurse, S1WantScold,
S1WantShout, S2WantCurse,
… S4DoShout.
As we expected, our Lord’s Wald test results varied when using the three item sets, because the item parameter estimates varied when using a different item set. When analyzing the targeted item “S1wantCurse” using each of the item sets A, B and C, we found Lord’s test statistic is 0.0006, 0.0002, 0.0246, and 1.3724, respectively. All these statistics are not statistically significant at the 0.05 level. They suggest that no DIF exists between the two gender groups treated as the reference group and the focal group in expressing the tendency to respond aggressively to a frustrating situation, concerning their answers to the survey item “S1wantCurse.” We used four item sets here only to demonstrate that the composition of the item set has an impact on the DIF results.
Table 8: Lord’s Wald test results.
Item $\chi^{2}_{Lord}$ $p$-value
A 0.0006 0.9800
B 0.0002 0.9899
C 0.0246 0.8753
D 1.3724 0.2414
  • ${}^{*}p<0.05$; ${}^{**}p<0.01$; ${}^{***}p<0.001$.
The following R code demonstrates the use of an existing R function difLord() to complete Lord’s Wald test for DIF analysis of the targeted item “S1wantCurse.” The full item set (composed by 24 items) are used for this analysis, and they are calibrated using the one-parameter item response model:
1library(difR)
2data("verbal")
3
4d4 = verbal[, c(1:24, 26)]
5res4 <- difLord(Data = d4, group = "Gender", focal.name = 1, model = "1PL")
6
7res4
8#// display the test statistic
Appendix C Selected Categorical Variables from the Codebook
Table 9: The item response variable “ADDPREV.”
Code Description
1 Yes
2 No
85 BAD DATA Logically assigned
94 DON’T KNOW
97 REFUSED
98 BLANK (NO ANSWER)
99 LEGITIMATE SKIP
Table 10: The group membership variable “PDEN10.”
Code Description
1 Segment in a CBSA with 1 million or more persons
2 Segment in a CBSA with fewer than 1 million persons
3 Segment not in a CBSA
Table 11: The level-2 variable “IREDUHIGHST2.”
Code Description
1 Fifth grade or less grade completed
2 Sixth grade completed
3 Seventh grade completed
4 Eighth grade completed
5 Ninth grade completed
6 Tenth grade completed
7 Eleventh or Twelfth grade completed, no diploma
8 High school diploma/GED
9 Some college credit, but no degree
10 Associate’s degree (for example, AA, AS)
11 College graduate or higher

### v2

@@ 4.1 Data preparation
The original NSDUH data file contains 2,605 variables and 59,069 respondents, composed by 27,047 males and 32,022 females aged 12 and older. We obtained two different data sets from this original data file. In the first data set, we reserve responses recoded with 1 for the response “Yes” and 0 for “No”, retain 45,827 eligible respondents to enable multilevel binary LR. In the second data set, we include responses recoded with 2 for the other valid responses including “Don’t know”, “Refused”, and blank answers, retaining 47,094 eligible respondents for multilevel multinominal LR. Both the data sets contain no missing values. Table 1 displays sample sizes by level-2 cluster.
Table 1: Sample size of respondents, by cluster.
Level-2 clusters Group Response
Edu. level (0) Low-density area (1) High-density area (2) Non-classified area (0) No (1) Yes (2) Others
Data Set 1: With binary responses
   1 126 140 10 232 44
   2 75 115 5 177 18
   3 45 49 7 82 19
   4 158 102 56 240 76
   5 314 206 39 414 145
   6 407 273 57 529 208
   7 1228 1059 142 1691 738
   8 6488 4430 746 7942 3722
   9 5208 3895 447 5607 3943
   10 2286 1563 226 2616 1459
   11 7630 7874 421 10029 5896
Data Set 2: With polytomous responses
   1 129 148 10 232 44 11
   2 77 120 5 177 18 7
   3 47 52 7 82 19 5
   4 165 104 58 240 76 11
   5 329 214 40 414 145 24
   6 422 289 57 529 208 31
   7 1270 1082 146 1691 738 69
   8 6716 4579 769 7942 3722 400
   9 5337 3985 456 5607 3943 228
   10 2350 1603 232 2616 1459 110
   11 7791 8075 430 10029 5896 371
  • Note. Numbers in parentheses in the header are codes in the data.
We keep the following four variables for DIF analysis:
@@
    The item response variable (named “ADDPREV” in the survey). It contains responses 0 and 1 for multilevel binary LR modeling, and it include the response 2 for multilevel multinominal LR modeling.
@@
    The group membership variable (named “PDEN10”). It indicates the population density of the area where respondents are residing. It is recoded to have 0 for “low-density” (i.e., with fewer than one million people), 1 for “high-density” (i.e., with one million or more people), and 2 for areas that are not involved in Core Based Statistical Area (CBSA) classifications. In modeling, this variable is dummy-coded.
@@
Appendix B displays information from the NSDUH codebook about how the selected categorical variables were originally coded. Appendix C.1 documents the R code for this data preparation.
@@ 4.2 Analysis and Results
The modeling results are in Table 2 and Table 3. The models in Table 2 are associated with the multilevel binary LR in (1) for analyzing binary responses, while the models in Table 3 are related to the multilevel multinominal LR in (2) for analyzing polytomous responses. The models in Table 2 were fitted via lme4 package [42] and the models in Table 3 were fitted via brms and rstan packages, with R4.4.0 in RStudio. Appendix C.2 documents the R code for this modeling.
In both the tables, the first three models are for multilevel model building, named “Model 0,” “Model 1,” and “Model 2,” and the last model is the single-level model, labeled with “Model 3.” Model 0 does not contain explanatory variables in the level-1 submodel, commonly known as null model or empty model that sets a baseline for understanding the proportion of variance explained by the hierarchical structure of the data [30]. Model 1 is the exact model used in (1) or (2), incorporating the latent trait and group membership variables. Model 2 adds the interaction term between the latent trait and group membership, which is the interaction term we drop from the traditional LR, to check how the model performance would change with this interaction.
Table 2: Multilevel binary LR: Analyzing binary responses.
@@
  N 45827 45827 45827 45827
  ICC 0.07 0.06 0.06
  AIC 59233 46286 46288 46809
$R^{2}$ 0.07 0.38 0.38
  • Note. ${}^{*}p<0.05$; ${}^{**}p<0.01$; ${}^{***}p<0.001$. $g_{p.1}$ and $g_{p.2}$ correspond to $g_{p}$ when the group membership is 1 and 2, respectively.
Table 3: Multilevel multinominal LR: Analyzing polytomous responses.
Model 0 Model 1 Model 2 Model 3
$Eq_{1}$ $Eq_{2}$ $Eq_{1}$ $Eq_{2}$ $Eq_{1}$ $Eq_{2}$ $Eq_{1}$ $Eq_{2}$
Fixed effects
  (Intercept) -1.02∗∗∗ -3.12∗∗∗ -2.37∗∗∗ -3.66∗∗∗ -2.37∗∗∗ -3.69∗∗∗ -1.89∗∗∗ -3.73∗∗∗
(0.19) (0.07) (0.18) (0.09) (0.18) (0.09) (0.02) (0.06)
$\theta_{p}$ 0.25∗∗∗ 0.13∗∗∗ 0.25∗∗∗ 0.14∗∗∗ 0.24∗∗∗ 0.14∗∗∗
(0.00) (0.01) (0.00) (0.01) (0.00) (0.01)
$g_{p.1}$0.12∗∗∗ -0.02 -0.09 0.03 -0.07 0.01
(0.02) (0.06) (0.04) (0.09) (0.04) (0.09)
$g_{p.2}$ 0.01 -0.14 -0.06 -0.17 -0.23 -0.29
(0.05) (0.15) (0.09) (0.22) (0.09) (0.22)
$\theta_{p}g_{p.1}$0.01 -0.01 0.00 -0.01
(0.01) (0.01) (0.01) (0.01)
$\theta_{p}g_{p.2}$ 0.01 0.01 0.04 0.06
(0.01) (0.03) (0.01) (0.03)
Level-2 Variances
$\tau_{0}^{2}$ 0.37 0.02 0.31 0.02 0.32 0.02
  N 47094 47094 47094 47094 47094 47094 47094 47094
  ICC 0.10 0.01 0.09 0.01 0.09 0.01
  WAIC 70824 70824 57909 57909 57915 57915 58506 58506
$R^{2}$ 0.00 0.00 0.23 0.23 0.23 0.23 0.23
  • Note. ${}^{*}p<0.05$; ${}^{**}p<0.01$; ${}^{***}p<0.001$. $g_{p.1}$ and $g_{p.2}$ correspond to $g_{p}$ when the focal group’s group membership value is 1 and 2, respectively. $Eq_{1}$ and $Eq_{2}$ refer to “Equation 1” and “Equation 2” when the targeted response is 1 and 2, respectively. They share the same WAIC and $R^{2}$ values for each model.
The modeling results reveal several key insights regarding the model performance. Model 0 serves as the baseline, showing the proportion of variance attributable to the hierarchical structure of the data. Its intraclass correlation coefficient [ICC; 25] is 0.07 in Table 2 and 0.10 in Table 3, indicating that 7% of the response variability stemmed from differences between education levels for the first data set and this proportion is 10% for the second data set. These values highlight that accounting for the hierarchical structure in DIF analyses matters with both the data sets.
Model 1 displays a significant negative effect of being from high-density areas on the probability of responding “Yes” to the item, for both the data sets. This suggests that the selected survey item has population-density-associated DIF, as individuals residing in high-density areas were less likely to provide the targeted response to the selected item compared to those in low-density areas, after accounting for all these individuals’ latent trait about depression. Model 3, the single-level model, does not reveal a significant effect of high population density, which leads to the conclusion that there is no DIF when comparing high- and low-density areas, contrasting with the result from Model 1. This difference underscores the importance of model selection when conducting DIF analyses.
Overall, Model 1 outperforms others, based on model performance statistics. Its intercept variance $\tau_{0}^{2}$ is among the lowest in both the tables, indicating an advantage of Model 1 in capturing more data variability. Additionally, Model 1 has the lowest value in the model fit index, namely the Akaike information criterion (AIC) in Table 2 and the widely applicable information criterion (WAIC) in Table 3. In addition, Model 1 has the highest conditional $R^{2}$ [43], signifying that the entire model, with random effects and fixed effects in combination, explains 38% of the variability in the first data set and 23% in the second data set.
@@ 5 Discussion
Our empirical analyses cover the DIF analysis of a dichotomous item and a polytomous item, using multilevel binary LR and multlevel multinominal LR in addition to their single-level counterparts. Rather than making a blanket recommendation in favor of multilevel models, we intend to raise awareness of how to use multilevel modeling to analyze DIF and how to choose the appropriate model when a series of competing models are available. Different from some other studies that evaluated models via Monte Carlo simulation [e.g., 44], our empirical analysis examines a series of competing models based on model performance statistics. Our findings indicate that multilevel modeling more effectively captures variability in our data compared to single-level modeling. Future research could explore the generalizability of our findings by applying multilevel DIF analysis to other self-reported health survey data. Also, researchers might consider comparing our models with other models capable of multilevel modeling for DIF analysis with polytomous items or more than two groups, ideally models without IRT-related assumptions such as ordinal LR [45, 46] or nonparametric measures [e.g., 47].
More research is needed to support multilevel modeling. First, literature limitations exist regarding model construction, ICC calculation, and generalizability [24], despite existing efforts to explore multilevel DIF analysis [e.g., 48, 49, 50]. Second, it would be worthwhile to examine the cost-benefit balance of multilevel modeling for DIF analysis in dynamic health survey settings, such as using Monte Carlo simulations to systematically evaluate model performance under a wide range of data structures. These simulations can vary factors such as ICC, number and size of clusters, the magnitude and proportion of DIF items, and data distribution within each cluster [30]. Studies by Moineddin et al. [51] and [28] provide valuable groundwork, but further exploration is necessary. Third, scrutinizing model performance under various realistic scenarios that mimic real-world health survey data is essential. This includes exploring the impact of factors like missing data patterns (missing completely at random, missing at random, or missing not at random) and different types of DIF (uniform vs. non-uniform) on model performance. By addressing these limitations, we can solidify the role of multilevel modeling in DIF analysis for health research, ultimately leading to more robust and reliable estimates captured through self-reported surveys.